General Meeting Summary 4/14/22

From Cuawiki
Jump to: navigation, search

PARTICIPANTS: Aaron Brown, Alexandre Camsonne, Charles Hyde, Arshak Asaturyan, Fernando Barbosa, Carlos Munoz-Camacho, Tanja Horn, Rolf Ent, Abishek Karki, Brad Sawatzky, Erik Wrighton, Hamlet Mkrtchyan, Steven Lassiter, Marie Boer, Vladimir Berdnikov, Paulo Medeiros

HV DIVIDERS

  • received ten first articles - tested - matched the reference
  • proceeded with production - some delays with Samtex connectors, but resolved now
  • in process of assembling - ~20% complete - expect to the on schedule


FRAME TEMPERATURE AND HUMIDITY SENSORS

  • DSG provided model - each sensor requires 4 slots in the keysight terminal block
  • Determine if enough space for cables etc. inside the enclosure
  • 24 (12x2) slots available in front 2 blocks
  • instead of placing humidity sensors inside on the frame attach them on the panel at front
  • at back no issue - enough space - need long cables though
  • Need to purchase: sensors, boards, cables
  • Action item: send drawings and any other information to Aaron
  • Also needed: modification of the top plate/panel
  • Brad can coordinate


UPDATES FROM DSG

  • EPICS HV Controls
  • Impact of cooling shell (10 degC) - thermal analysis with no continous heat load (at ambient temperature) - 210 crystals in central zone hottest (between 19.0 and 20.5 degC)
  • concentrated at back - impacts only tail of shower
  • need to review longitudinal heat map
  • any impact on resolution?
  • still need to include cooling at back
  • assume 0.5 W/channel,
  • Carbon fiber thermal conductivity
  • .5523w/m*degc
  • back mu metal has higher conductivity (19 W/m*K)
  • Is there an option to turn on/off HV row by row (or column by column)
  • working on turning off all channels for one module
  • Air flow at back of the crystals, no flow at front part of crystal
  • Model air as fluid? - not yet
  • Action items
  • collect all information in repository
  • evaluate if the material could be published in a journal


MECHANICAL/ENGINEERING

  • NPS deck has been installed almost completely
  • Stairs in place - needs to be mounted
  • Note: it's a temporary fitting - needs to be taken off again for rotation to reach 5 deg angle in PionLT


INSTALLATION SCHEDULE

  • Revisited to assist with scheduling the next experiments in Hall C
  • 41 days from beam off (1 March 2023) until all support structures ready (HB removal, take down beam line, platform install and fitting, cable tray layout, cable connections, etc.) - total is ~2-3 months
  • Roof block - presently consider modifying existing roof block
  • Mechanically can be ready by 1 May 2023 - assume until 29 May 2023 for cable hookup and checkout
  • Experiment schedule - will NPS experiment be scheduled or need to go for Jeopardy


ASSEMBLY PREPARATIONS

  • Assembly space - on track to be available at end of April
  • Shelving - request was put in, but not yet available - alternatives available
  • Action item:
  • Brad and Vladimir will further discuss offline
  • Onsite personnel - in particular for May as this is where crystal wrapping has to be completed before the IJCLab-Orsay team arrives for stacking
  • for wrapping ideally need two shifts of ~4 hours each per day with each shift 2-3 people
  • so far the following teams have indicated availability: AANL (2), CUA (2-3), DSG (1-2 at the end of May only)
  • close, but still room for additional collaborators - will follow up by email
  • Action items:
  • collect number of collaborators on site in May
  • organize a shift schedule for the crystal wrapping
  • Safety documents preparation - Brad will check
  • DAQ/Electronics and Analyzer
  • moving forward
  • meeting every 1-2 weeks organized by David Hamilton


NEXT MEETING: THURSDAY MAY 12 AT 9:00AM ET